# A Game-Theoretic Unital Classification Theorem for $C^{*}$-Algebras

### v1

@@ 2.4. Refining families of Borel equivalence relations
We consider in this section analytic sets $X$ and $Y$ in some respective ambient standard Borel spaces, together with equivalence relations $E\subseteq X^{2}$ and $F\subseteq Y^{2}$. Suppose we are given a map $K\colon X\to Y$ such that $x\ E\ x^{\prime}$ holds if and only if $K(x)\ F\ K(x^{\prime})$. Suppose further that $F$ is the intersection of a decreasing sequence $(F_{\alpha})_{\alpha<\omega_{1}}$ of equivalence relations on $Y$. It is natural to ask whether classification already occurs at lower levels of the sequence of equivalence relations, i.e., whether there is some $\beta<\omega_{1}$ such that $x\ E\ x^{\prime}$ already holds whenever $K(x)\ F_{\beta}\ K(x^{\prime})$. One can ask a similar question in the case where $E$ is the intersection of such a sequence. In general there is no reason this should be the case, but we show in this section that under suitable descriptive set theoretic assumptions on X,Y,E,FX,~Y,~E,~F and $K$ such $\beta<\omega_{1}$ exists.
We use the following classical result by Lusin and Sierpiński (see [MR1321597, Theorem 27.12]). Denote by $\operatorname{WO}\subseteq\operatorname{LO}$ the set of well-orders on subsets of $\omega$. Note that $\operatorname{LO}$ is Borel and $\operatorname{WO}$ is co-analytic.
@@
By Suslin’s theorem, non-analyticity of $\operatorname{WO}$ is a consequence of non-Borelness of $\operatorname{WO}$, or equivalently (by the Lopez-Escobar theorem) non-definability of $\operatorname{WO}$ in $L_{\omega_{1},\omega}$. This can be seen as a rather weak fraction of compactness that is preserved in infinitary logic: For $L=\{<,\dots\}$ a language and $\sigma\in L_{\omega_{1},\omega}$ a sentence, if $\sigma$ has models of order-type $\alpha$ for all $\alpha<\omega_{1}$, then $\sigma$ has a model of the order-type of $(\mathbb{Q},<)$.
@@
Call a family $R\subseteq X\times\mathcal{I}$ indexed by $\mathcal{I}\subseteq\operatorname{LO}$ refining, if $R_{I}\supseteq S_{I^{\prime}}$ whenever $I,I^{\prime}\in\mathcal{I}$ are such that II embeds into $I^{\prime}$.
Denote by $\omega^{*}$ the linear order obtained by inverting the usual order on $\omega$. As a consequence of Lusin-Sierpiński, we obtain the following,
@@
Let $R\subseteq X\times\operatorname{LO}$ be analytic and refining. If for every countable ordinal $\alpha<\omega_{1}$, the set $R_{\alpha}$ is non-empty, then $R_{\omega^{*}}$ is non-empty.
@@
By our assumption this set contains all well-orders on $\omega$ and it does not contain $\omega^{*}$. Since any non-well-order contains a copy of $\omega^{*}$ and $R$ is refining, we conclude that $R_{I}=\emptyset$ for all non-well-orders $I$. Thus, $\mathcal{I}=\operatorname{WO}$. Moreover, $\mathcal{I}$ is analytic as a projection of the analytic set $R$, which contradicts Theorem 2.22. ∎
@@
A family $E=(E_{\alpha})_{\alpha<\omega_{1}}\in\mathcal{P}(X^{2})\times\omega_{1}$ is called a refining Borel (analytic) equivalence relation on $X$ if $E_{\alpha}\subseteq X^{2}$ is a Borel (analytic) equivalence relation for each $\alpha<\omega_{1}$ and $E_{\alpha}\supseteq E_{\beta}$ whenever $\alpha<\beta<\omega_{1}$. $E$ is called uniformly Borel (analytic) if there is a Borel (analytic) refining ${R}\subseteq X^{2}\times\operatorname{LO}$ with $E_{\alpha}=R_{\alpha}$ for each $\alpha<\omega_{1}$ and $R_{\omega^{*}}=\bigcap_{\alpha}E_{\alpha}$. In this case, we say $R$ induces $E$.
@@
Then $R^{\prime}$ is analytic as a projection of an analytic set, since $R$, $K$, and $X^{2}\setminus E$ are analytic. Moreover, $R^{\prime}$ is refining and $R^{\prime}_{\beta}\neq\emptyset$ for each $\beta<\omega_{1}$, by assumption and the fact that $F_{\beta}\subseteq R_{\beta}$. By Lemma 2.24, $R^{\prime}_{\omega^{*}}$ is non-empty. Thus, there are $x,x^{\prime}\in X$ such that $\neg x\ E\ x^{\prime}$ and $(Kx,Kx^{\prime})\in R_{\omega^{*}}$. However, $R_{\omega^{*}}\subset F_{\infty}$, so $K(x)\ F_{\infty}\ K(x^{\prime})$ and thus $x\ E_{\infty}\ x^{\prime}$ by assumption. In particular $x\ E\ x^{\prime}$, which gives a contradiction. ∎
Note that for the above proof, it would have been sufficient to assume co-analyticity of $E$ in place of Borelness. A converse of Theorem 2.27 can be proved analogously, assuming uniform analyticity of the sequence of equivalence relations $(E_{\alpha})_{\alpha<\omega_{1}}$, and that $K$ maps $E_{\infty}$-equivalent pairs to $F_{\infty}$-equivalent pairs.
@@
Again, note that if would have been sufficient to assume co-analyticity of $F$, in place of Borelness.
We turn now to the case where $X$ and $Y$ are spaces of separable metric structures. Let $L_{1}$ and $L_{2}$ be metric languages. Recall that $\equiv_{\alpha}$ denotes elementary equivalence up to quantifier-rank $\alpha$ and $\equiv_{\alpha}^{\text{bf}}$ is back-and-forth equivalence with respect to a fixed universal weak modulus in the respective language, in the sense of Definition 2.10. We may write ${\equiv}^{L_{i}}_{\alpha}$ and ${\equiv}_{\alpha}^{L_{i},\text{bf}}$ to highlight the dependence on the ambient language. As an immediate application of the above theorem and the work of Section 2.3, we obtain Corollary C. In the following statement (and its proof) we identify $A\in\operatorname{Mod}_{\omega}(L)$ with the separable $L$-structure $\widehat{A}$ it encodes.
@@
$K(A)\cong K(B)\iff A\cong B.$
@@
$\begin{split}K(A)\equiv_{\theta(\alpha)}K(B)\implies A\equiv_{\alpha}B,\\ A\equiv_{\theta^{\prime}(\alpha)}B\implies K(A)\equiv_{\alpha}K(B).\end{split}$
@@
Note that for $i\in\{0,1\}$, both $({\equiv}^{L_{i}}_{\alpha})_{\alpha<\omega_{1}}$ and $\equiv_{\alpha}^{L_{i},bf}$ are refining sequences of equivalence relations, with limit $\cong$, and moreover $(\equiv_{\alpha}^{L_{2},bf})_{\alpha<\omega_{1}}$ is uniformly Borel by Theorem 2.21. Applying Theorem 2.27 with $E=E_{\alpha}$ and Theorem 2.28 with $F=F_{\alpha}$, for each $\alpha<\omega_{1}$, yields functions $\theta\colon\omega_{1}\to\omega_{1}$ and $\theta^{\prime}\colon\omega_{1}\to\omega_{1}$ such that for any $A,B\in\operatorname{Mod}_{\omega}(L_{1})$ and $\alpha<\omega_{1}$,
@@ 2.5. Borel category of $L$-structures
We introduce now the notion of a Borel category, and construct a Borel model for the category of separable $C^{*}$-algebras. More generally, we construct a Borel model of the category of separable models of any given theory in a countable metric language. If one considers only surjective isometries between models, such Borel groupoids were already considered in [ElliottEtAl2013, MR3660238] (using the Urysohn space) and in [yaacov2017metric]. Characterizations of Borel functors between Borel groupoids of models were given in [MR3893282, MR4029718]. Let us note that the study of Borel structures (see [steinhorn1985chapter, MR2830412] for the definition) was initiated by Friedman [friedman1979borel] and there are many results knows for some ‘degenerated’ Borel categories, like orderings or equivalence relations [MR965754, MR1057041].
@@
Let $M$ and $N$ be $L$-structures for some metric language $L$. A homomorphism $M\to N$ is a map $\psi\colon M\to N$ such that $\psi$ commutes with interpretations of function and constant symbols in $L$ and for any relation symbol $R(x)$ in $L$ and $a\in M^{x}$,
@@
In particular, taking $R=d$ the distinguished metric symbol ensures that homomorphisms are contractive.
For an $L_{\omega_{1},\omega}$-sentence $\sigma$ (or a countable collection of such), write $\mathcal{C}_{\sigma}$ for the category of separable models of $\sigma$, with homomorphisms as morphisms.
@@
We fix a countable collection of $L_{\omega_{1},\omega}$-sentences $T$ for some metric language $L$. By Lemma 2.8, the set $\operatorname{Mod}_{\omega}(T)\subset\operatorname{Mod}_{\omega}(L)$ is Borel. The goal now is to construct a Borel category $\mathbf{X}^{T}$ with $X^{T}_{\operatorname{Ob}}:=\operatorname{Mod}_{\omega}(T)$, and to find a functor $F\colon\mathbf{X}^{T}\to\mathcal{C}_{T}$ witnessing equivalence of categories. We use the convention that if we talk about functors from Borel categories to usual categories we mean functors from the category underlying the Borel one. For $A\in X_{\operatorname{Ob}}$, we will denote by $\widehat{A}$ the separable $L$-structure that $A$ encodes, i.e., $\widehat{A}$ is the set of equivalence classes of Cauchy sequences in $A$, with functions and relations extended continuously (so until the end of this section we do not identify $A$ with $\widehat{A}$). At the level of objects, it is clear that $F$ will coincide with $\widehat{\cdot}$, which we have already seen to be essentially surjective (i.e., every separable model can be encoded in $\operatorname{Mod}_{\omega}(T)$). We are left with finding a suitable space of morphisms in $\mathbf{X}^{T}$. For $A,B\in\operatorname{Mod}_{\omega}(T)$, we want to encode homomorphisms $\widehat{f}\colon\widehat{A}\to\widehat{B}$ in a way that only depends on $A$ and $B$. We do this in the natural way, using Cauchy sequences.
@@
Let $X_{1}\subset X^{T}_{\operatorname{Ob}}\times X^{T}_{\operatorname{Ob}}\times(\omega^{\omega})^{\omega}$ be the set of triples $(A,B,f)$ with $f=(f_{n})_{n\in\omega}$ satisfying the following conditions.
@@
    For every $a\in\omega$, the sequence $f_{n}(a)$ is Cauchy with respect to $d_{B}$;
@@
    There is a (unique) homomorphism $\widehat{f}\colon\widehat{A}\to\widehat{B}$ such that $\widehat{f}(a)=\lim_{n}f_{n}(a)\in\widehat{B}$ for each $a\in\omega$.
@@
    $\bigwedge_{a\in A}\bigwedge_{\varepsilon>0}\bigvee_{N}\bigwedge_{n,m>N}d_{B}(f_{n}(a),f_{m}(a))<\varepsilon$
    which is clearly Borel as we only used countable intersections and unions and the condition $d_{B}(f_{n}(a),f_{m}(a))<\varepsilon$ is open.
@@
    Let $R$ be some relation symbol of arity $k$. We check that respecting $R$ in the sense of (3) is a Borel condition. The relevant formula in this case is the following:
    $\bigwedge_{a\in A^{k}}\bigwedge_{\varepsilon>0}\bigvee_{N\in\omega}\bigwedge_{n\geq N}|R^{B}(f_{n}(a))|<|R^{A}(a)|+\varepsilon,$
    which is clearly Borel. Similarly, for $\rho$ a $k$-ary function symbol in $L$, commuting with $\rho$ is a Borel condition, defined by
    $\bigwedge_{a\in A^{k}}\bigwedge_{\varepsilon>0}\bigvee_{N\in\omega}\bigwedge_{n\geq N}d_{B}\big(f_{n}(\rho^{A}(a)),\rho^{B}(f_{n}(a))\big)<\varepsilon.$
    Note that we had to use the fact that $A$ is closed under $g^{A}$, to make sense of the expression $g^{A}(a)$.
Note that the above conditions are not just necessary, but also sufficient for $(A,B,(f_{n})_{n\in\omega})\in X_{1}$. Indeed, any $(f_{n})_{n\in\omega}$ with these properties induces a map $f^{\prime}\colon A\to\widehat{B}$ with $f^{\prime}(a):=\lim_{n}f_{n}(a)$, and this map is continuous with respect to $d_{A}$ and $d_{B}$, as can be seen by applying $(ii)$ to $R=d$. It follows that $f^{\prime}$ can uniquely be extended to a continuous $\widehat{f}\colon\widehat{A}\to\widehat{B}$, and this will be a homomorphism by (ii) and density of $A$. ∎
@@
$A=C,~B=D\wedge\bigwedge_{a\in A}\bigwedge_{b\in B}\bigwedge_{\varepsilon>0}\bigvee_{N\in\omega}\bigwedge_{n\geq N}\big|d_{f}(a,b)-d_{B}(f_{n}(a),b)\big|<\varepsilon,$
which is clearly a Borel relation. The second assertion follows, since $\lim_{n}f_{n}(a)=\lim_{n}g_{n}(a)$ in $\widehat{B}$ if and only if $d_{B}(\lim f_{n}(a),b)=d_{B}(\lim g_{n}(a),b)$ for all $b\in B$, by density of $B$ in $\widehat{B}$.
It is clear now that $\pi(X_{1})$ is analytic, since $X_{1}$ is Borel by Lemma 2.36 and $\pi$ is a Borel map. By Suslin’s theorem (Fact 2.1), it is enough to check that $\pi(X_{1})$ is co-analytic, which we do now.
Let $(A,B,\rho)\in X^{T}_{\operatorname{Ob}}\times X^{T}_{\operatorname{Ob}}\times{\mathbb{R}}^{\omega\times\omega}$. Then $(A,B,\rho)\notin\pi(X_{1})$ if and only if $(A,B,\rho)$ satisfies one of the following conditions.
  • $h$ does not correspond to a function. The first case is that there are Cauchy sequences $(a_{n})$ in $A$, $(b_{n}),(c_{n})$ in $B$ with $\lim_{n}b_{n}\neq\lim_{n}c_{n}$ in $\widehat{B}$ and
    $\bigwedge_{\varepsilon>0}\bigvee_{N\in\omega}\bigwedge_{n\geq N}\max(\rho(a_{n},b_{n}),\rho(a_{n},c_{n}))<\varepsilon.$
    The second case is that there is a Cauchy sequence $a_{n}$ that does not map to any element in $B$, i.e.,
    $\bigvee_{\varepsilon>0}\bigwedge_{b\in B}\bigwedge_{n}\rho(a_{n},b)>\varepsilon;$
    Note that both of these expressions are Borel conditions in $(A,B,\rho)$ and $(a_{n}),(b_{n}),(c_{n})$. The set of Cauchy sequence with respect to fixed $A\in X^{T}_{\operatorname{Ob}}$, respectively $B\in X^{T}_{\operatorname{Ob}}$, is Borel in $\omega^{\omega}$ [^7] This can be seen via the same proof as in Lemma 2.36., and similarly the condition that $\lim_{n}b_{n}\neq\lim_{n}c_{n}$ in $\widehat{B}$ is Borel . Thus, the set of $(A,B,\rho)$ satisfying one of the above cases is analytic, as the projection (coming from existential quantification over the Cauchy sequences) of a Borel set.
  • $\rho$ corresponds to a function that does not induce a homomorphism of L{L}-structures as in Definition 2.34. That is, there is some symbol in $L$ that $\rho$ does not respect. Take, for example, a binary relation $R$. The relevant condition on $h$ then is that there are $a,a^{\prime}\in A$ and Cauchy sequences $(b_{n}),(b_{n}^{\prime})$ in $B$, with
    $\bigwedge_{\varepsilon>0}\bigvee_{N\in\omega}\bigwedge_{n\geq N}\max(\rho(a,b_{n}),\rho(a^{\prime},b_{n}^{\prime}))<\varepsilon,$
    and
    $\bigvee_{\varepsilon>0}\bigwedge_{N}\bigvee_{n>N}|R(b_{n},b^{\prime}_{n})|\geq|R(a,a^{\prime})|+\varepsilon.$
    The case of a function symbol is similar. Again, this set is analytic as a projection of a Borel set, and we take the countable union over all symbols in $L$ of such sets, which is still analytic.

### v2

@@ 2.4. Refining families of Borel equivalence relations
Consider sets $X$ and $Y$ together with equivalence relations $E\subseteq X^{2}$ and $F\subseteq Y^{2}$. Suppose we are given a map $K\colon X\to Y$ such that $x\ E\ x^{\prime}$ holds if and only if $K(x)\ F\ K(x^{\prime})$. Suppose further that $F$ is the intersection of a decreasing sequence $(F_{\alpha})_{\alpha<\omega_{1}}$ of equivalence relations on $Y$. It is natural to ask whether classification already occurs at lower levels of the sequence of equivalence relations, i.e., whether there is some $\beta<\omega_{1}$ such that $x\ E\ x^{\prime}$ already holds whenever $K(x)\ F_{\beta}\ K(x^{\prime})$. One can ask a similar question in the case where $E$ is the intersection of such a sequence. In general there is no reason this should be the case, but we show in this section that under suitable descriptive set theoretic assumptions on X,Y,E,FX,~Y,~E,~F and $K$ such $\beta<\omega_{1}$ exists.
We use the following classical result by Lusin and Sierpiński (see [39, Theorem 27.12]). Denote by $\operatorname{WO}\subseteq\operatorname{LO}$ the set of well-orders on subsets of $\omega$. Recall that $\operatorname{LO}$ denotes the set of linear orders on subsets of $\omega$. Observe that $\operatorname{LO}$ is Borel and $\operatorname{WO}$ is co-analytic.
@@
By Suslin’s theorem, non-analyticity of $\operatorname{WO}$ is a consequence of non-Borelness of $\operatorname{WO}$, or equivalently (by the Lopez-Escobar theorem) non-definability of $\operatorname{WO}$ in $L_{\omega_{1},\omega}$ whenever $L=\{<,\dots\}$. This can be seen as a rather weak fraction of compactness that is preserved in infinitary logic: If $L=\{<,\dots\}$ and $\sigma\in L_{\omega_{1},\omega}$ is a sentence, if $\sigma$ has models of order-type $\alpha$ for all $\alpha<\omega_{1}$, then $\sigma$ has a model of the order-type of $(\mathbb{Q},<)$.
@@
Call a family $R\subseteq X\times\mathcal{I}$ indexed by $\mathcal{I}\subseteq\operatorname{LO}$ refining, if $R_{I}\supseteq R_{I^{\prime}}$ whenever $I,I^{\prime}\in\mathcal{I}$ are such that II embeds into $I^{\prime}$.
Denote by $\omega^{*}$ the linear order obtained by inverting the usual order on $\omega$. Denote by $X$ and $Y$ analytic subsets of respective standard Borel spaces. As a consequence of Lusin-Sierpiński, we obtain the following,
@@
Let $R\subseteq X\times\operatorname{LO}$ be analytic and refining. If $R_{\alpha}\neq\emptyset$ for every $\alpha<\omega_{1}$, then $R_{\omega^{*}}\neq\emptyset$.
@@
By assumption this set contains all well-orders on $\omega$ and it does not contain $\omega^{*}$. Since any non-well-order contains a copy of $\omega^{*}$ and $R$ is refining, we conclude that $R_{I}=\emptyset$ for all non-well-orders $I$. Thus, $\mathcal{I}=\operatorname{WO}$. Moreover, $\mathcal{I}$ is analytic as a projection of the analytic set $R$, contradicting Theorem 2.22. ∎
@@
A refining family $E=(E_{\alpha})_{\alpha<\omega_{1}}\in\mathcal{P}(X^{2})\times\omega_{1}$ is called a refining Borel (analytic) equivalence relation on $X$ if $E_{\alpha}\subseteq X^{2}$ is a Borel (analytic) equivalence relation for each $\alpha<\omega_{1}$. $E$ is called uniformly Borel (analytic) if there is a Borel (analytic) refining ${R}\subseteq X^{2}\times\operatorname{LO}$ with $E_{\alpha}=R_{\alpha}$ for each $\alpha<\omega_{1}$ and $R_{\omega^{*}}=\bigcap_{\alpha<\omega_{1}}E_{\alpha}$. In this case, we say $R$ induces $E$.
@@
Then $R^{\prime}$ is analytic as a projection of an analytic set, since $R$, $K$, and $\neg E$ are analytic. Moreover, $R^{\prime}$ is refining and $R^{\prime}_{\beta}\neq\emptyset$ for each $\beta<\omega_{1}$, by assumption and the fact that $F_{\beta}=R_{\beta}$. By Lemma 2.24, $R^{\prime}_{\omega^{*}}$ is non-empty. Thus, there are $x,x^{\prime}\in X$ such that $\neg x\ E\ x^{\prime}$ and $(Kx,Kx^{\prime})\in R_{\omega^{*}}$. However, $R_{\omega^{*}}=F_{\infty}$, so $K(x)\ F_{\infty}\ K(x^{\prime})$ and thus $x\ E\ x^{\prime}$ by assumption. ∎
Note that for the above proof, it would have been sufficient to assume co-analyticity of $E$ in place of Borelness. Moreover, it would be enough to have $F_{\beta}\subseteq R_{\beta}$ for each $\beta<\omega_{1}$ and $R_{\omega^{*}}\subseteq F_{\infty}$. A converse of Theorem 2.27 can be proved analogously, assuming uniform analyticity of the sequence of equivalence relations $(E_{\alpha})_{\alpha<\omega_{1}}$, and that $K$ maps $E_{\infty}$-equivalent pairs to $F$-equivalent pairs.
@@
Again, note that it is sufficient to assume co-analyticity of $F$, in place of Borelness.
We turn now to the case where $X$ and $Y$ are spaces of separable metric structures. Let $L_{1}$ and $L_{2}$ be metric languages. Recall that $\equiv_{\alpha}$ denotes elementary equivalence up to quantifier-rank $\alpha$ and $\equiv_{\alpha}^{\text{bf}}$ is back-and-forth equivalence in the sense of Definition 2.10 with respect to a fixed universal weak modulus in the respective language. We may write ${\equiv}^{L_{i}}_{\alpha}$ and ${\equiv}_{\alpha}^{L_{i},\text{bf}}$ to highlight the dependence on the ambient language. As an immediate application of the above theorem and Example 2.26, we obtain the following, amounting to Corollary C. In the following statement (and its proof) we identify $A\in\operatorname{Mod}_{\omega}(L)$ with the separable $L$-structure $\widehat{A}$ it encodes.
@@
$\forall A,B\in X\colon\ K(A)\cong K(B)\iff A\cong B.$
@@
$\forall A,B\in X\ \forall\alpha<\omega_{1}\colon\begin{cases}K(A)\equiv_{\theta(\alpha)}K(B)\implies A\equiv_{\alpha}B,\\ A\equiv_{\theta^{\prime}(\alpha)}B\implies K(A)\equiv_{\alpha}K(B).\end{cases}$
@@
Note that for $i\in\{0,1\}$, both $({\equiv}^{L_{i}}_{\alpha})_{\alpha<\omega_{1}}$ and $\equiv_{\alpha}^{L_{i},bf}$ are refining sequences of equivalence relations, and $\bigcap_{\alpha}{\equiv}^{L_{i}}_{\alpha}$ and $\bigcap_{\alpha}\equiv_{\alpha}^{L_{i},bf}$ coincide with $\cong$. Moreover, $(\equiv_{\alpha}^{L_{2},bf})_{\alpha<\omega_{1}}$ is uniformly Borel by Theorem 2.21. Applying Theorem 2.27 with $E=E_{\alpha}$ and Theorem 2.28 with $F=F_{\alpha}$, for each $\alpha<\omega_{1}$, yields functions $\theta\colon\omega_{1}\to\omega_{1}$ and $\theta^{\prime}\colon\omega_{1}\to\omega_{1}$ such that for any $A,B\in\operatorname{Mod}_{\omega}(L_{1})$ and $\alpha<\omega_{1}$,
@@ 2.5. Borel category of $L$-structures
We introduce now the notion of a Borel category, and construct a Borel model for the category of separable $C^{*}$-algebras. More generally, we construct a Borel model of the category of separable models of any given theory in a countable metric language. If one considers only surjective isometries between models, such Borel groupoids were already considered in [12, 44] (using the Urysohn space) and in [2]. Characterizations of Borel functors between Borel groupoids of models were given in [28, 7]. Let us note that the study of Borel structures (see [51, 34] for the definition) was initiated by Friedman [19] and there are many results knows for some ‘degenerated’ Borel categories, like orderings or equivalence relations [27, 26].
@@
Fix some metric language $L$. We aim now to find a Borel model of the category of separable $L$-structures, with homomorphisms in the sense of the following definition as morphisms.
@@
Let $M$ and $N$ be $L$-structures. A homomorphism $M\to N$ is a map $\psi\colon M\to N$ such that $\psi$ commutes with interpretations of function and constant symbols in $L$ and for any relation symbol $R(x)$ in $L$ and $a\in M^{x}$,
@@
For an $L_{\omega_{1},\omega}$-theory $T$, write $\mathcal{C}_{T}$ for the category of separable models of $T$, with homomorphisms as morphisms.
Note that in particular all homomorphisms are contractive, by (2) with $R$ being the distinguished symbol $d$.
@@
We fix a countable collection of $L_{\omega_{1},\omega}$-sentences $T$ for some metric language $L$. By Lemma 2.8 and the fact that $T$ is countable, the set $\operatorname{Mod}_{\omega}(T)\subset\operatorname{Mod}_{\omega}(L)$ is Borel. The goal now is to construct a Borel model $\mathbf{X}^{T}$ of $\mathcal{C}_{T}$ with $X^{T}_{\operatorname{Ob}}:=\operatorname{Mod}_{\omega}(T)$, i.e. to find a functor $F\colon\mathbf{X}^{T}\to\mathcal{C}_{T}$ witnessing equivalence of categories. We use the convention that if we talk about functors from Borel categories to usual categories we mean functors from the category underlying the Borel one. For $A\in X_{\operatorname{Ob}}$, we will denote by $\widehat{A}$ the separable $L$-structure that $A$ encodes, i.e., $\widehat{A}$ is the set of equivalence classes of Cauchy sequences in $A$, with functions and relations extended continuously (so until the end of this section we do not identify $A$ with $\widehat{A}$). At the level of objects, it is clear that $F$ will coincide with $\widehat{\cdot}$, which we have already seen to be essentially surjective (i.e., every separable model can be encoded in $\operatorname{Mod}_{\omega}(T)$). We are left with finding a suitable space of morphisms in $\mathbf{X}^{T}$. For $A,B\in\operatorname{Mod}_{\omega}(T)$, we want to encode homomorphisms $\widehat{f}\colon\widehat{A}\to\widehat{B}$ in a way that only depends on $A$ and $B$. We do this in the natural way, using Cauchy sequences.
@@
Let $X_{1}\subset X^{T}_{\operatorname{Ob}}\times X^{T}_{\operatorname{Ob}}\times(\omega^{\omega})^{\omega}$ be the set of triples $(A,B,f)$ satisfying the following conditions.
@@
    For every $a\in A$, the sequence $f_{n}(a)$ is Cauchy with respect to $d_{B}$;
@@
    There is a (unique) homomorphism $\widehat{f}\colon\widehat{A}\to\widehat{B}$ such that $\widehat{f}(a)=\lim_{n}f_{n}(a)\in\widehat{B}$ for each $a\in A$.
@@
    $\bigwedge_{a\in A}\bigwedge_{\varepsilon>0}\bigvee_{N<\omega}\bigwedge_{n,m>N}d_{B}(f_{n}(a),f_{m}(a))<\varepsilon$
    which is Borel as we only used countable intersections and unions and the condition $d_{B}(f_{n}(a),f_{m}(a))<\varepsilon$ is open.
@@
    Let $R$ be some relation symbol of arity $k$. We check that $\widehat{f}$ respecting $R$ in the sense of (2) is a Borel condition on $(A,B,f)$. The condition is captured by the formula
    $\bigwedge_{a\in A^{k}}\bigwedge_{\varepsilon>0}\bigvee_{N<\omega}\bigwedge_{n\geq N}|R^{B}(f_{n}(a))|<|R^{A}(a)|+\varepsilon,$
    which is clearly Borel. Note it is crucial here that $R^{\widehat{B}}$ is continuous, i.e. that $R^{\widehat{B}}(\widehat{f}(a))=\lim_{n}R^{B}(f_{n}(a))$. Similarly, for a $k$-ary function symbol $\rho$, the property that $\widehat{f}$ commutes with $\rho$ is defined by
    $\bigwedge_{a\in A^{k}}\bigwedge_{\varepsilon>0}\bigvee_{N<\omega}\bigwedge_{n\geq N}d_{B}\big(f_{n}(\rho^{A}(a)),\rho^{B}(f_{n}(a))\big)<\varepsilon,$
    which is Borel in $(A,B,f)$. Note that we again used continuity of $\rho^{\widehat{B}}$; and the fact that $A$ is closed under $\rho^{A}$ for well-definedness of $f_{n}(\rho^{A}(a))$.
The above conditions are not just necessary, but also sufficient for $(A,B,(f_{n})_{n\in\omega})\in X_{1}$. Indeed, any $(f_{n})_{n\in\omega}$ with these properties induces a map $f^{\prime}\colon A\to\widehat{B}$ with $f^{\prime}(a):=\lim_{n}f_{n}(a)$, and this map is continuous (in fact contractive) with respect to $d_{A}$ and $d_{B}$, as can be seen by applying $(ii)$ to $R=d$. It follows that $f^{\prime}$ can uniquely be extended to a continuous $\widehat{f}\colon\widehat{A}\to\widehat{B}$, and this will be a homomorphism by (ii) and density of $A$. ∎
@@
$A=B~\wedge~B=D~\wedge~\bigwedge_{a\in A}\bigwedge_{b\in B}\bigwedge_{\varepsilon>0}\bigvee_{N<\omega}\bigwedge_{n\geq N}\big|d_{f}(a,b)-d_{B}(f_{n}(a),b)\big|<\varepsilon,$
which is a Borel condition. The second assertion follows, since $\lim_{n}f_{n}(a)=\lim_{n}g_{n}(a)$ in $\widehat{B}$ if and only if $d_{B}(\lim f_{n}(a),b)=d_{B}(\lim g_{n}(a),b)$ for all $b\in B$, by density of $B$ in $\widehat{B}$.
It follows also that $\pi(X_{1})$ is analytic, since $X_{1}$ is Borel by Lemma 2.36 and $\pi$ is a Borel map. By Suslin’s theorem (Fact 2.1), we are left with checking that $\pi(X_{1})$ is co-analytic.
Let $(A,B,\delta)\in X^{T}_{\operatorname{Ob}}\times X^{T}_{\operatorname{Ob}}\times{\mathbb{R}}^{\omega\times\omega}$. Then $(A,B,\delta)\notin\pi(X_{1})$ if and only if $(A,B,\delta)$ satisfies one of the following conditions.
  • $\delta$ does not induce a function. The first case is that there is $a\in A$, and Cauchy sequences $(b_{n}),(c_{n})\in B^{\omega}$ w.r.t. $d_{B}$ with $\lim_{n}b_{n}\neq\lim_{n}c_{n}$ and $\lim_{n}\delta(a,b_{n})=\lim_{n}\delta(a,c_{n})=0$. The second case is that there is some $a\in A$ that does not map to any element in $\widehat{B}$, that is,
    $\bigvee_{\epsilon>0}\bigwedge_{b\in B}\delta(a,b)>\epsilon.$
    Note that both of these expressions are Borel conditions in $(A,B,\rho)$ and $a,(b_{n}),(c_{n})$. The set of Cauchy sequence with respect to $B$ is Borel in $\omega^{\omega}$ [^7] This can be seen via the same proof as in Lemma 2.36., and similarly the condition that $\lim_{n}b_{n}\neq\lim_{n}c_{n}$ in $\widehat{B}$ is Borel. Thus, the set of triples $(A,B,\delta)$ satisfying one of these two cases is analytic, as the projection (coming from existential quantification over the Cauchy sequences) of a Borel set.
  • $\delta$ corresponds to a function that does not induce a homomorphism of L{L}-structures as in Definition 2.34. Consider, for example, a binary relation $R$. Then the function induced by $\delta$ fails to satisfy (2) if there are $a,a^{\prime}\in A$ and Cauchy sequences $(b_{n})$ and $(b_{n}^{\prime})$ in $B$ with $\lim_{n}\delta(a,b_{n})=\lim_{n}\delta(a^{\prime},b_{n}^{\prime})=0$ and
    $\bigvee_{\varepsilon>0}\bigwedge_{N<\omega}\bigvee_{n>N}|R(b_{n},b^{\prime}_{n})|\geq|R(a,a^{\prime})|+\varepsilon.$
    Again, this condition is analytic as a projection of a Borel set. The case of a function symbol or constant symbol is similar. We take the countable union over all symbols in $L$ of these conditions, which is still analytic.